OpenStudy (anonymous):

whats the discriminant for F(x)=6x^2+5x-3

OpenStudy (kirbykirby):

\(ax^2+bx+c\), here: a=6 b=5 c=-3 \(discriminant = b^2-4ac = 5^2-4(6)(-3)\)